OpenStudy (anonymous):

Determine all primes \(p\) for which the system \(p + 1 = 2x^2\) \(p^2 + 1 = 2y^2\) has a solution in integers \(x\); \(y\).

OpenStudy (anonymous):

sure u did this... \(p(p-1)=2(y-x)(y+x)\)

OpenStudy (anonymous):

and use this \(1<x<y<p\) since without loss of generality we can assume \(x\) , \(y\) are positive integers

OpenStudy (experimentx):

since -1 is not prime ... i guess the answer would only be 7

OpenStudy (anonymous):

how u found out that \(p=x+y\)

OpenStudy (experimentx):

p(p-1) = 2(x+y)(x-y) since p ! = 2*anything and p > p - 1 p-1 = 2(x-y) and p = x+y

OpenStudy (anonymous):

\(p\) is odd so \(p | (y-x)(y+x) \) we know that \(p>y-x\) hence \(p | x+y\)....on the other hand \(x+y< 2p\) so \(p=x+y\)